Calculate delta Hvap for water from your plot (m= -delta Hvap / R)
Carry out the calculations fo the guron .tor only one tennperotura volume measudomextr Find moles aur at 5 C. L-0,002L. air atm COID k mal
media%2F483%2F48346c3b-3ace-442b-bccf-96
Answer #1

nAir = 759.2 Torr * (0.0038L – 0.002L) / 62.4 * 5°C + 279 K)

        = 759.2 * 0.0018 / 62.4 * 284

        = 1.366 / 17721.599 = 7.708 * 10^-5 moles of air.

If the plot of ln P vs 1/T is made slope (m) = -?Hvap/R

   1 Tinc TinV(L, corrected) n(air) P(water) In P 1/T air 75 348.15 70 343.15 65 338.15 60 333.15 55 328.15 50 323.15 5 278.15 2